OpenStudy (anonymous):

(x-3)(x+2)=0 Do you factor to solve for 0?

OpenStudy (anonymous):

now it is already factored

OpenStudy (anonymous):

I got x^2 +2x-3x-6=0

OpenStudy (anonymous):

what -3 is 0 and what + 2 is 0 are your 2 answers

OpenStudy (anonymous):

use the zero product property ssince product is zero, one of the factors has to be zero if x - 3 = 0, then x = 3 if x + 2 = 0, then x = -2

OpenStudy (anonymous):

zero product property! i rememeber that, ty Harkirat!

OpenStudy (anonymous):